Please solve by substitution or elimination.
-3x-3y=-15
8x+6y=24

Answer by Matthewcus11(2)

Simultaneous Equation by elimination method
-3x-3y=-15
8x+6y=24
(-3x-3y=-15)X6
(8x+6y=24)X-3
-18x-18y=-90
-24x-18y=-72
-
6x =-18
/6
x=-3
let equation 2 be use for solving y
8x+6y =24
8(-3)+6y=24
-24+6y=24
6y =24+24=48
/6
6y=48
/6 /6
y=8